OpenStudy (anonymous):

Solve the quadratic equation by completing the square. 3x2 + x - 5 = 0

OpenStudy (anonymous):

you meant this, \[3x^2 + x - 5 = 0\] (Not a perfect example to complete the square)

OpenStudy (anonymous):

yes and i know its hard

OpenStudy (anonymous):

Well, it would be easier if you multiply everything times 3, this way _x^2 will be a perfect square at least.

OpenStudy (anonymous):

do you know the answer?

OpenStudy (anonymous):

\[3x^2 + x - 5 = 0~~~~~~~~~~~~~\times~3\]\[9x^2 + 3x - 15 = 0\]\[9x^2 + 3x = 15\]\[(3x+1/2)(3x+1/2 )=15\]now you do it.

OpenStudy (anonymous):

finish it please im writing a paper to

OpenStudy (anonymous):

\[(3x+1/2)=\sqrt{15}\]\[3(x+1/6)=\sqrt{15}\]\[x+1/6=+-\sqrt{15}/3\]\[x=-1/6+-\sqrt{15}/3\]\[x=-1/6+-2\sqrt{15}/6\]\[x=(~-1+-2\sqrt{15}~)~/~6\]